Quinine is a natural product extracted from the bark of the cinchona tree, which is native to South America. Quinine is used as an antimalarial agent. When 0.950 g of quinine is dissolved in 25.0 g of cyclohexane, the freezing point of the solution is lowered by 2.44 ∘ C. Look up the freezing point and K f constant for cyclohexane in the Colligative Constants table. Calculate the molar mass of quinine.
Answer #1

Kf = 20.20C/m

Tf = 2.440C

i = 1 for non electrolyte

Tf   = i*Kf*m

2.44   = 1*20.2*m

m = 2.44/20.2 = 0.12m

molality   = W*1000/G.M.Wt * weight of solvent in g

0.12         = 0.95*1000/G.M.Wt *25

G.M.Wt   = 0.95*1000/(0.12*25)   = 316.67g/mole

The molar mass of quinine.   = 316.67g/mole >>>>>answer